# Convolution estimates for the Boltzmann gain operator with hard-spheres

### v3

@@ 4.4. The $L_{v}^{r}$– estimate for $Q_{1}$
${\langle}v{\rangle}^{k}|Q_{1}^{s_{2}}(f,g)|$ $\lesssim{\langle}v{\rangle}^{1/r}\|f_{l}\|_{L_{v}^{p_{1}}}^{p_{1}/q_{1}^{\prime}}\|g_{k}\|_{L_{v}^{q_{1}}}^{q_{1}/p^{\prime}_{1}}$
(4.50) $\quad\times\left(\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|f_{l}(v^{*})|^{p_{1}}|g_{k}(v_{1}^{*})|^{q_{1}}\mathds{1}_{|v|>10}\mathds{1}_{(v_{1},\sigma)\in\mathcal{B}_{2}(v)}\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1}\right)^{1/r}.$
Raising the above inequality to the $r$-th power and integrating in $v$, we obtain
$\|{\langle}v{\rangle}^{k}Q_{1}^{s_{2}}(f,g)\|_{L_{v}^{r}}^{r}$ $\lesssim\|f_{l}\|_{L_{v}^{p_{1}}}^{r-p_{1}}\|g_{k}\|_{L_{v}^{q_{1}}}^{r-q_{1}}$
(4.51) $\quad\times\left(\int_{{\mathbb{R}}^{6}\times{\mathbb{S}}^{2}}{\langle}v{\rangle}|f_{l}(v^{*})|^{p_{1}}|g_{k}(v_{1}^{*})|^{q_{1}}\mathds{1}_{|v|>10}\mathds{1}_{(v_{1},\sigma)\in\mathcal{B}_{2}(v)}\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1}\,dv\right)$
Recalling (4.16), and using (4.41) we obtain
(4.52) $\mathds{1}_{|v|>10}\mathds{1}_{(v_{1},\sigma)\in\mathcal{B}_{1}(v)}\chi(\widehat{u}\cdot\sigma)\leq\mathds{1}_{E>100}\mathds{1}_{|v|>\frac{E^{1/2}}{6}}\mathds{1}_{|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2}<6E^{-1/2}}\chi(\widehat{u}\cdot\sigma).$
Using (4.52), followed by (2.12) and (2.28), the integral term of (4.51) can be estimated as
$\int_{{\mathbb{R}}^{6}\times{\mathbb{S}}^{2}}$ ${\langle}v{\rangle}|f_{l}(v^{*})|^{p_{1}}|g_{k}(v_{1}^{*})|^{q_{1}}\mathds{1}_{|v|>10}\mathds{1}_{(v_{1},\sigma)\in\mathcal{B}_{2}(v)}\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1}\,dv$
$\leq\int_{{\mathbb{R}}^{6}\times{\mathbb{S}}^{2}}E^{1/2}|f_{l}(v^{*})|^{p_{1}}|g_{k}(v_{1}^{*})|^{q_{1}}\mathds{1}_{E>100}\mathds{1}_{|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2}<6E^{-1/2}}\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1}\,dv$
$=\int_{{\mathbb{R}}^{6}\times{\mathbb{S}}^{2}}E^{1/2}|f_{l}(v^{*})|^{p}|g_{k}(v_{1}^{*})|^{q}\mathds{1}_{E>100}\mathds{1}_{\widehat{u}\cdot\sigma>1-12E^{-1/2}}\,d\sigma\,dv_{1}\,dv$
$=\int_{{\mathbb{R}}^{6}}E^{1/2}|f_{l}(v)|^{p_{1}}|g_{k}(v_{1})|^{q_{1}}\mathds{1}_{E>100}\int_{{\mathbb{S}}^{2}}\mathds{1}_{\widehat{u}\cdot\sigma>1-12E^{-1/2}}\,d\sigma\,dv_{1}\,dv$
$=\int_{{\mathbb{R}}^{6}}E^{1/2}|f_{l}(v)|^{p_{1}}|g_{k}(v_{1})|^{q_{1}}\mathds{1}_{E>100}\int_{1-12E^{-1/2}}^{1}\,dx\,dv_{1}\,dv$
(4.53) $\approx\int_{{\mathbb{R}}^{6}}|f_{l}(v)|^{p_{1}}|g_{k}(v_{1})|^{q_{1}}\,dv_{1}\,dv=\|f_{l}\|_{L_{v}^{p_{1}}}^{p_{1}}\|g_{k}\|_{L_{v}^{q_{1}}}^{q_{1}}.$
By (4.51), (4.53), we conclude
(4.54) $\|{\langle}v{\rangle}^{k}Q_{1}^{s_{2}}(f,g)\|_{L_{v}^{r}}\lesssim\|f_{l}\|_{L_{v}^{p_{1}}}\|g_{k}\|_{L_{v}^{q_{1}}}.$
@@
We conclude with the proof of Theorem 1.1. Using decomposition (4.1), the triangle inequality and Propositions 4.14.4, we obtain
(4.55) $\|{\langle}v{\rangle}^{k}Q^{+}(f,g)\|_{L_{v}^{r}}\lesssim\|{\langle}v{\rangle}^{k}f\|_{L_{v}^{p}}\|{\langle}v{\rangle}^{l}g\|_{L_{v}^{q}}+\|{\langle}v{\rangle}^{l}f\|_{L_{v}^{p}}\|{\langle}v{\rangle}^{k}g\|_{L_{v}^{q}}+\|{\langle}v{\rangle}^{l}f\|_{L_{v}^{q}}\|{\langle}v{\rangle}^{k}g\|_{L_{v}^{p}},$
for all $f,g\in C_{c}^{\infty}({\mathbb{R}}^{3})$. Recalling the definition of $\|\cdot\|_{\bm{\mathcal{X}_{p,q}^{k}}}$ from (1.19), the above bound implies
(4.56) $\|{\langle}v{\rangle}^{k}Q^{+}(f,g)\|_{L_{v}^{r}}\lesssim\|f\|_{\bm{\mathcal{X}_{p,q}^{k}}}\|g\|_{\bm{\mathcal{X}_{p,q}^{k}}},\quad\forall f,g\in C_{c}^{\infty}({\mathbb{R}}^{3}).$
By Lemma A.1, we can uniquely extend $Q^{+}$ by density to a bounded bilinear operator $Q^{+}:\bm{\mathcal{X}_{p,q}^{k}}\times\bm{\mathcal{X}_{p,q}^{k}}\to{\langle}v{\rangle}^{-k}L^{r}_{v}$ by a standard density argument.
Finally, we prove that (1.21) holds for any $f,g\in\bm{\mathcal{X}_{p,q}^{k}}$. Indeed, consider $f,g\in\bm{\mathcal{X}_{p,q}^{k}}$. Using Lemma A.1, we pick sequences $\{f_{n}\}_{n=1}^{\infty}$, $\{g_{n}\}_{n=1}^{\infty}\subset C_{c}^{\infty}({\mathbb{R}}^{3})$ with $f_{n}\to f$ and $g_{n}\to g$ as $n\to\infty$ in $\bm{\mathcal{X}_{p,q}^{k}}$. Then, by (4.55), for any $n\in\mathbb{N}$ we have
$\|{\langle}v{\rangle}^{k}Q^{+}(f_{n},g_{n})\|_{L_{v}^{r}}\lesssim\|{\langle}v{\rangle}^{k}f_{n}\|_{L_{v}^{p}}\|{\langle}v{\rangle}^{l}g_{n}\|_{L_{v}^{q}}+\|{\langle}v{\rangle}^{l}f_{n}\|_{L_{v}^{p}}\|{\langle}v{\rangle}^{k}g_{n}\|_{L_{v}^{q}}+\|{\langle}v{\rangle}^{l}f_{n}\|_{L_{v}^{q}}\|{\langle}v{\rangle}^{k}g_{n}\|_{L_{v}^{p}}.$
Letting $n\to\infty$ in the above bound, and using continuity of $Q^{+}$ for the left-had side as well as (1.20) for the right-hand side, estimate (1.21) follows. The proof is complete.

### v4

@@ 4.2. The $L_{v}^{r}$– estimate
$\int_{0}^{1}\frac{\mathds{1}_{\sqrt{1-x^{2}}<{\langle}v{\rangle}^{-1/2}}}{(1-x^{2})^{p/q}}\,dx\lesssim\int_{0}^{1}\frac{\mathds{1}_{1-x<{\langle}v{\rangle}^{-1}}}{(1-x)^{q^{\prime}/p^{\prime}}}\,dx\approx\int_{0}^{{\langle}v{\rangle}^{-1}}z^{-q^{\prime}/p^{\prime}}\,dz\lesssim{\langle}v{\rangle}^{\frac{q^{\prime}}{p^{\prime}}-1},$
where we used the fact that $p<q$ for the convergence of the integral in $z$. We obtain
$J_{1}\lesssim\int_{{\mathbb{R}}^{3}}|\phi(v)|^{r^{\prime}}\int_{{\mathbb{R}}^{3}}|f_{l}(v-\nu)|^{p}\,d\nu\,dv=\|f_{l}\|_{L^{p}_{v}}^{p}\|\phi\|_{L^{r^{\prime}}_{v}}^{r^{\prime}}.$
To estimate $J_{3}$, we note that in the corresponding domain of integration we have $|v|\gtrsim E^{1/2}$. Indeed, by (2.12) we have $|\widehat{R}_{\sigma}^{-}\cdot\sigma|<{\langle}v{\rangle}^{-1/2}<1/2$, since $|v|>4$. Now, by (2.11), (1.6), we have
$|R_{\sigma}^{-}(u)|=|u||\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|<\frac{|u|}{2}=\frac{|v^{*}-v_{1}^{*}|}{2}\leq\frac{|v_{1}^{*}|}{2}+\frac{|v^{*}|}{2}\leq\frac{|v_{1}^{*}|}{2}+\frac{|v|}{4},$
since $|v^{*}|<|v|/2$. Then, using (2.7) and the triangle inequality, we obtain
$|v|\geq|v_{1}^{*}|-|R_{\sigma}^{-}(u)|\geq\frac{|v_{1}^{*}|}{2}-\frac{|v|}{4}\Rightarrow|v|\geq\frac{2|v_{1}^{*}|}{5}>\frac{2E^{1/2}}{5\sqrt{2}}.$
Therefore, using the above estimate, the fact that $|v|\leq E^{1/2}$, and (2.12), (2.28), we can bound
$J_{3}$ $\lesssim\int_{{\mathbb{R}}^{6}\times{\mathbb{S}}^{2}}E^{1/2}|f_{l}(v^{*})|^{p}|g_{k}(v_{1}^{*})|^{q}\mathds{1}_{E>16}\mathds{1}_{|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2}<E^{-1/2}}\,d\sigma\,dv_{1}\,dv$
$=\int_{{\mathbb{R}}^{6}\times{\mathbb{S}}^{2}}E^{1/2}|f_{l}(v^{*})|^{p}|g_{k}(v_{1}^{*})|^{q}\mathds{1}_{E>16}\mathds{1}_{\widehat{u}\cdot\sigma>1-2E^{-1/2}}\,d\sigma\,dv_{1}\,dv$
$=\int_{{\mathbb{R}}^{6}}E^{1/2}|f_{l}(v)|^{p}|g_{k}(v_{1})|^{q}\mathds{1}_{E>16}\int_{{\mathbb{S}}^{2}}\mathds{1}_{\widehat{u}\cdot\sigma>1-2E^{-1/2}}\,d\sigma\,dv_{1}\,dv$
$\approx\int_{{\mathbb{R}}^{6}}E^{1/2}|f_{l}(v)|^{p}|g_{k}(v_{1})|^{q}\mathds{1}_{E>16}\int_{1-2E^{-1/2}}^{1}\,dx\,dv_{1}\,dv$
$\approx\int_{{\mathbb{R}}^{6}}|f_{l}(v)|^{p}|g_{k}(v_{1})|^{q}\,dv_{1}\,dv=\|f_{l}\|_{L_{v}^{p}}^{p}\|g_{k}\|_{L_{v}^{q}}^{q}.$
Combining (4.18) with the bounds on $J_{1},J_{2},J_{3}$, we deduce
$|J(\phi)|\lesssim\|f_{l}\|_{L^{p}_{v}}\|g_{k}\|_{L^{q}_{v}}\|\phi\|_{L^{r^{\prime}}_{v}}.$
Since $\phi\in L^{r^{\prime}}_{v}$ was chosen arbitrarily, we conclude
(4.19) $\|{\langle}v{\rangle}^{k+1-\gamma}Q_{s,2}^{+}(f,g)\|_{L_{v}^{r}}\lesssim\|f_{l}\|_{L_{v}^{p}}\|g_{k}\|_{L_{v}^{q}}.$
@@
We conclude with the proof of Theorem 1.1. Using Propositions 4.14.2, we obtain
(4.20) $\|{\langle}v{\rangle}^{k+1-\gamma}Q^{+}(f,g)\|_{L_{v}^{r}}\lesssim\|{\langle}v{\rangle}^{l}f\|_{L_{v}^{p}}\|{\langle}v{\rangle}^{k}g\|_{L_{v}^{q}}+\|{\langle}v{\rangle}^{l}f\|_{L_{v}^{q}}\|{\langle}v{\rangle}^{k}g\|_{L_{v}^{p}},\quad\forall\,f,g\in C_{c}({\mathbb{R}}^{3}).$
Recalling the definition of $\|\cdot\|_{\bm{\mathcal{X}_{p,q}^{k}}}$ from (1.17), the above bound implies
(4.21) $\|{\langle}v{\rangle}^{k+1-\gamma}Q^{+}(f,g)\|_{L_{v}^{r}}\lesssim\|f\|_{\bm{\mathcal{X}_{p,q}^{k}}}\|g\|_{\bm{\mathcal{X}_{p,q}^{k}}},\quad\forall f,g\in C_{c}({\mathbb{R}}^{3}).$
By Lemma A.1, we can uniquely extend $Q^{+}$ by density to a bounded bilinear operator $Q^{+}:\bm{\mathcal{X}_{p,q}^{k}}\times\bm{\mathcal{X}_{p,q}^{k}}\to{\langle}v{\rangle}^{-k-1+\gamma}L^{r}_{v}.$
Finally, we prove that (1.19) holds for any $f,g\in\bm{\mathcal{X}_{p,q}^{k}}$. Indeed, consider $f,g\in\bm{\mathcal{X}_{p,q}^{k}}$. Using Lemma A.1, we choose sequences $\{f_{n}\}_{n=1}^{\infty}$, $\{g_{n}\}_{n=1}^{\infty}\subset C_{c}^{\infty}({\mathbb{R}}^{3})$ such that $f_{n}\to f$ and $g_{n}\to g$ as $n\to\infty$ in $\bm{\mathcal{X}_{p,q}^{k}}$. Then, by (4.20), for any $n\in\mathbb{N}$ we have
$\|{\langle}v{\rangle}^{k+1-\gamma}Q^{+}(f_{n},g_{n})\|_{L_{v}^{r}}\lesssim\|{\langle}v{\rangle}^{l}f_{n}\|_{L_{v}^{p}}\|{\langle}v{\rangle}^{k}g_{n}\|_{L_{v}^{q}}+\|{\langle}v{\rangle}^{l}f_{n}\|_{L_{v}^{q}}\|{\langle}v{\rangle}^{k}g_{n}\|_{L_{v}^{p}}.$
Letting $n\to\infty$ in the above bound, and using continuity of $Q^{+}$ for the left-hand side as well as (1.18) for the right-hand side, estimate (1.19) follows. The proof is complete.